OpenStudy (anonymous):

Which of the following points is a solution of y ≤ -|x| - 1? a)(-1, -3) b)(0, 0) c)(1, -1)

OpenStudy (anonymous):

(a). would be the only correct answer... If you substitute the x=-1 you will get y\[\le\]-|-1|-1... thus y\[\le\]-1-1 ... y\[\le\]-2... since -3 is less than or equal to -2, (a) is the only correct answer from the choices given.

OpenStudy (anonymous):

Answer: A when you plug in the value of X given in the point list. You will see that (-1,-3) in the one which fit well. y <= -|-1| -1 then y <= -1 -1 .....y<= -2 so -3 is less than -2 therefore the answer is (-1,-3)